# FlatCAD: Fast Curvature Regularization of Neural SDFs for CAD Models

### v1

@@ 3.2. Loss Function Components
Our training objective combines four terms that constrain the signed‐distance field (SDF) both on and off the surface. The total loss is
(3) $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{DM}}+\lambda_{\text{DNM}}\mathcal{L}_{\text{DNM}}+\lambda_{\text{eik}}\mathcal{L}_{\text{eikonal}}+\lambda_{\text{proxy}}\mathcal{L}_{\text{proxy}},$
with scalar weights $\lambda_{\text{DNM}}$, $\lambda_{\text{eik}}$, and $\lambda_{\text{proxy}}$ balancing the contributions of the three regularizers against the data term.
Manifold Constraint
For the $N$ input points $\{\mathbf{x}_{i}\}_{i=1}^{N}$ sampled on or very near the target surface, we impose a Dirichlet condition that pulls their SDF values to 00:
@@
(6) $\mathcal{L}_{\text{eikonal}}=\frac{1}{K}\sum_{k=1}^{K}\Bigl{(}\,\lVert\nabla f(\mathbf{z}_{k})\rVert_{2}-1\Bigr{)}^{2}.$
This global term regularizes gradient norms, stabilizing optimization and improving extrapolation away from the data.
Curvature Proxy (ours)
@@
(7) $\mathcal{L}_{\text{proxy}}=\frac{1}{L}\sum_{\ell=1}^{L}\bigl{|}S_{12}(\theta_{\ell})\bigr{|}.$
The mixed second derivative $S_{12}$ is evaluated with a single Hessian–vector product, avoiding explicit Hessian assembly.
@@
Weights. We keep most parameter weights identical to the settings recommended in NeurCADRecon: $\lambda_{\text{DM}}=7000$, $\lambda_{\text{DNM}}=600$, $\lambda_{\text{eik}}=50$, and $\lambda_{\text{proxy}}=10$.
@@ 4. Computational Pathways
The mixed–curvature term $S_{12}$ can be evaluated with or without explicit second-order autodifferentiation, giving two practical routes that trade accuracy for running cost. The finite-difference minimizes memory use, whereas the autodiff variant offers higher accuracy. Both achieve the curvature-gap regularization without the overhead of full second-order methods.
4.1. First-Order Proxy via Taylor Expansion
When raw throughput or framework simplicity is paramount we approximate the mixed derivative with the central stencil of Equation (1). This requires four additional forward evaluations of the SDF at offsets $h\mathbf{u}$, $h\mathbf{v}$, and $h(\mathbf{u}+\mathbf{v})$ plus a first-order gradient to normalise by $\|\nabla f\|$. Because no second-order graph is constructed, back-propagation resembles that of the standard Eikonal term; memory consumption is unchanged and the stencil error vanishes linearly with the step size $h$.
Evaluating the Hessian $Hf$ through automatic differentiation incurs full second-order cost. To reduce computation we derive a finite-difference estimator for the mixed second derivative $\mathbf{u}^{\!\top}Hf\,\mathbf{v}$, which is the only quantity required by our curvature-gap proxy.
Finite-difference stencil
Let $x_{\Omega}=x+d\,\nabla f/\|\nabla f\|$ be a point in the off–surface shell, and let
$f_{00}$ $=f(x_{\Omega}),$ $f_{\mathbf{u}}$ $=f(x_{\Omega}+h\,\mathbf{u}),$
$f_{\mathbf{v}}$ $=f(x_{\Omega}+h\,\mathbf{v}),$ $f_{\mathbf{uv}}$ $=f(x_{\Omega}+h\,\mathbf{u}+h\,\mathbf{v}),$
where $\mathbf{u},\mathbf{v}$ form an orthonormal tangent frame at $x_{\Omega}$ and $h>0$ is a small step size. The mixed central difference
(8) $D_{\mathbf{uv}}^{(+)}(x_{\Omega})=\frac{f_{\mathbf{uv}}-f_{\mathbf{u}}-f_{\mathbf{v}}+f_{00}}{h^{2}}$
approximates the desired Hessian entry.
Second-order expansion
Applying a second-order Taylor series at $x_{\Omega}$ gives
$f_{\mathbf{uv}}$ $=f_{00}+h\,\nabla f^{\!\top}(\mathbf{u}+\mathbf{v})+\tfrac{1}{2}h^{2}(\mathbf{u}+\mathbf{v})^{\!\top}Hf\,(\mathbf{u}+\mathbf{v})+\mathcal{O}(h^{3}),$
$f_{\mathbf{u}}$ $=f_{00}+h\,\nabla f^{\!\top}\mathbf{u}+\tfrac{1}{2}h^{2}\mathbf{u}^{\!\top}Hf\,\mathbf{u}+\mathcal{O}(h^{3}),$
$f_{\mathbf{v}}$ $=f_{00}+h\,\nabla f^{\!\top}\mathbf{v}+\tfrac{1}{2}h^{2}\mathbf{v}^{\!\top}Hf\,\mathbf{v}+\mathcal{O}(h^{3}).$
Substituting in $D_{\mathbf{uv}}^{(+)}$ yields
$D_{\mathbf{uv}}^{(+)}(x_{\Omega})\;=\;\mathbf{u}^{\!\top}Hf(x_{\Omega})\,\mathbf{v}\;+\;\mathcal{O}(h).$
Proxy evaluation.
@@ 4.1. First-Order Proxy via Taylor Expansion
$S_{12}(x_{\Omega})=\frac{\mathbf{u}^{\!\top}Hf(x_{\Omega})\,\mathbf{v}}{\|\nabla f(x_{\Omega})\|}.$
Using the finite-difference estimate gives the first-order proxy
(9) $|S_{12}(x_{\Omega})|\;\approx\;\frac{|D_{\mathbf{uv}}^{(+)}(x_{\Omega})|}{\|\nabla f(x_{\Omega})\|}\;+\;\mathcal{O}(h),$
@@
(10) $S_{12}^{2}(x_{\Omega})\;\approx\;\frac{\bigl{(}D_{\mathbf{uv}}^{(+)}(x_{\Omega})\bigr{)}^{2}}{\|\nabla f(x_{\Omega})\|^{2}}\;+\;\mathcal{O}(h).$
Rationale
The stencil converts the curvature-gap proxy into a purely first-order computation: four forward evaluations of $f$ replace one reverse-mode Hessian–vector product, eliminating second-order autodiff while maintaining an $\mathcal{O}(h)$ truncation error. Note that the expression involves only the mixed term $S_{12}$; it does not estimate the full Gaussian curvature $\det S$, which depends on all four components of the shape operator.
4.2. Second-Order Mixed Term
Autodiff Proxy
For applications that favor numerical precision we exploit reverse-mode autodiff to obtain a Hessian–vector product
@@ 4.2. Second-Order Mixed Term
The first backward sweep delivers $\nabla f$; a second sweep differentiates the scalar inner product $(\nabla f)\!\cdot\!\mathbf{v}$ with respect to the input coordinates, giving the exact mixed second derivative without materializing the full $3{\times}3$ Hessian. This doubles the backward cost for the proxy samples yet still keeps the overall computational graph first-order and avoids the prohibitive memory footprint of full Hessian assembly.
This Hessian–vector product therefore requires only two reverse-mode sweeps and stores no matrix entries, so its cost grows linearly with the number of shell samples and is independent of the ambient dimension. Dividing by $\|\nabla f(x_{\Omega})\|$ provides the exact $S_{12}(x_{\Omega})$ value used in the curvature-gap loss. Although more expensive than the first-order finite-difference proxy, this second-order route offers higher accuracy and serves as our reference implementation for ablation studies.
@@ 5. Experiments
In this section, we first describe the implementation details and evaluation metrics used in our validation. We then present a comprehensive comparison of our proposed proxy‐based loss functions (Proxy‐AD and Proxy‐FD) against three baseline methods—DiGS (Ben-Shabat et al., 2021), NSH (Wang et al., 2023), and the state‐of‐the‐art NeurCADRecon  (Dong et al., 2024)—on two subsets of the ABC dataset (Koch et al., 2019). Finally, we analyze convergence behavior and runtime performance.
@@ 5.1. Experimental Setup
All experiments are conducted on two 100‐example subsets of the ABC dataset (Koch et al., 2019). The first subset (1 MB set) is a pseudo‐random selection of 100 CAD models (each $\approx$ 1 MB in file size), while the second subset (5 MB set) comprises 100 human‐chosen models (between 3.5 MB and 9.5 MB, in average $\approx$ 5 MB) selected for their clean topology and well‐defined features. For each mesh, we uniformly sample a total of 30,000 points to form the input point cloud. This preprocessing strategy ensures that all methods receive identical input distributions. After that, in each iteration 20,000 points are randomly chosen from the 30,000 points. Additional 20,000 non-manifold points are drawn via spatial sampling within the mesh’s bounding volume.
@@
We compare two variants of our proxy loss—Proxy-AD (Auto-Diff) and Proxy-FD (Finite-Diff)—against three baselines: DiGS(Ben-Shabat et al., 2021), Neural Singular Hessian(Wang et al., 2023) and NeurCADRecon (Dong et al., 2024). All methods have been implemented within a unified comparison framework to eliminate discrepancies arising from differing network architectures or training pipelines. Each method employs a SIREN MLP architecture (Sitzmann et al., 2020) with four hidden layers of 256 units and sine activations. All models use a standard SIREN initialization, except DiGS, which uses the MFGI initialization. Training is performed using the Adam optimizer (Kingma and Ba, 2017) with a learning rate of $5\times 10^{-5}$ for up to 10,000 iterations. Early stopping is triggered if the Chamfer-Distance metric plateaus for 1500 consecutive iterations. All other hyperparameters, such as weight coefficients for curvature or normal losses, and other method specific parameters follow the settings from the original implementations.
@@ 5.2. Evaluation Metrics
We evaluate reconstruction accuracy using three standard metrics. Chamfer Distance (CD), scaled by $10^{3}$, measures the similarity between two surfaces; lower values indicate better accuracy. F1 Score (F1) is the harmonic mean of precision and recall, computed using a distance threshold of $5\times 10^{-3}$ between the predicted and ground-truth point sets. The results are scaled by $10^{2}$, with higher values indicating a better overlap. Normal Consistency (NC) captures the mean cosine similarity between predicted and ground-truth normals, scaled by $10^{2}$; higher values denote better alignment. For each of these metrics, we report the mean and standard deviation across all 100 shapes in each data subset. Additionally, for the 1MB subset, we include additional efficiency measures: GPU VRAM usage (in GB), Time per iteration (in milliseconds), and Best Iteration achieved during evaluation, the iteration count at which the best CD metric is reached. We utilize the time measurements and best iteration count to calculate the convergence time.
@@ 5.3. Qualitative Results
Beyond aggregate scores, the spatial distribution of reconstruction error is inspected in Figure 3 on a Hausdorff-distance heat map. Our method markedly reduce peak deviations and suppress local error spikes compared with the baselines. Other qualitative results are shown in Figure 1 and in the collection in Figure 10. We also compare our method with other state-of-the-art reconstructive methods, shown in Figure 4. Result shows our approach achieves comparable reconstruction quality with NeurCADRecon (Dong et al., 2024).
@@ 5.4. Input Data Experiments
Input Size NC $\mathrm{CD}_{L_{1}}$ F1
@@
To evaluate robustness to missing regions, a circular patch on the surface is excised from each input cloud while keeping the remaining samples unchanged. As reported in Table 2 and Figure 6, the proxy-based reconstruction degrades gracefully: Chamfer distance increases by only $64\,\%$ and normal consistency drops by $0.7\,\%$, yet global topology and salient features are preserved, however, a visible bump remains.
@@
Our method is specifically designed for CAD data, where the proxy loss encourages the reconstructed surface to be locally developable. To understand how this inductive bias affects performance on general, non-developable geometry, we evaluate our model on the standard Armadillo shape, which contains rich organic details and non-zero Gaussian curvature across much of its surface.
As shown in Figure 9 and Table 3, our method produces smooth and topologically consistent reconstructions, comparable to those generated by NeurCADRecon (Dong et al., 2024), but at approximately half the training time. Compared to DiGS (Ben-Shabat et al., 2021) and Neural-Singular-Hessian (NSH) (Wang et al., 2023), both our method and NeurCADRecon yield superior global shape recovery, exhibiting fewer holes and more coherent topology. However, we observe that fine surface details are partially smoothed out—an expected consequence of the developability regularization, which favors larger planar or gently curved regions.
@@
Figure 7. We compare the reconstructive shape with different proxy weight. From our result, our method consistently produces smooth and structurally faithful reconstructions across a wide range of weights.

### v2

@@ 3.4. Loss Function Components
Our training objective combines four terms that constrain the signed‐distance field both on and off the surface. The total loss is
(4) $\mathcal{L}_{\text{total}}=\lambda_{\text{DM}}\mathcal{L}_{\text{DM}}+\lambda_{\text{DNM}}\mathcal{L}_{\text{DNM}}+\lambda_{\text{EIK}}\mathcal{L}_{\text{EIK}}+\lambda_{\text{ODW}}\mathcal{L}_{\text{ODW}},$
with scalar weights $\lambda_{\text{DM}}$, $\lambda_{\text{DNM}}$, $\lambda_{\text{EIK}}$, and $\lambda_{\text{ODW}}$ balancing the contributions of the three regularizers against the data term.
Manifold Loss
For the $N$ input points $\{\mathbf{x}_{i}\}_{i=1}^{N}$ sampled on or very near the target surface, we impose a Dirichlet condition that pulls their SDF values to 0:
@@
(7) $\mathcal{L}_{\text{EIK}}=\frac{1}{K}\sum_{k=1}^{K}\Bigl{(}\,\lVert\nabla f(\mathbf{z}_{k})\rVert_{2}-1\Bigr{)}^{2}.$
This global term regularizes gradient norms, stabilizing optimization and improving extrapolation away from the data (Gropp et al., 2020).
Off-Diagonal Weingarten Loss (ours)
@@
(8) $\mathcal{L}_{\text{ODW}}=\frac{1}{L}\sum_{\ell=1}^{L}\bigl{|}S_{12}(\theta_{\ell})\bigr{|}.$
The mixed second derivative $S_{12}$ is evaluated either with a single Hessian–vector product or via the symmetric finite-difference approximation, avoiding explicit Hessian assembly.
@@
Weights. We keep most parameter weights identical to the settings recommended in NeurCADRecon: $\lambda_{\text{DM}}=7000$, $\lambda_{\text{DNM}}=600$, $\lambda_{\text{EIK}}=50$, and $\lambda_{\text{ODW}}=10$.
@@ 4. Computational Pathways
The mixed–curvature term $S_{12}$ can be evaluated with or without explicit second-order auto-differentiation, giving two practical routes that trade accuracy for running cost. The finite-difference minimizes memory use, whereas the auto-diff variant offers higher accuracy. Both achieve the curvature-gap regularization without assembling the full Hessian.
4.1. Off-Diagonal Weingarten Loss via Taylor Expansion
When raw throughput or framework simplicity is paramount we approximate the mixed derivative with a symmetric finite-difference stencil. This stencil evaluates the SDF at forward and backward offsets of magnitude $h$, averages the two estimates, and therefore incurs a truncation error of order $h^{2}$. If only the one-sided (forward) stencil is used, the truncation error is $\mathcal{O}(h)$ and the estimate is biased. Our default uses the symmetric average which cancels the odd powers and yields $\mathcal{O}(h^{2})$ accuracy.
Let $x_{\Omega}$ be any point in the near-surface shell $\Omega$, and let $\mathbf{u},\mathbf{v}$ form an orthonormal tangent frame at $x_{\Omega}$. For a small step $h>0$ define the forward offsets
$f_{00}$ $=f(x_{\Omega}),$ $\quad f_{\mathbf{u}}^{+}$ $=f(x_{\Omega}+h\,\mathbf{u}),$
$f_{\mathbf{v}}^{+}$ $=f(x_{\Omega}+h\,\mathbf{v}),$ $\quad f_{\mathbf{uv}}^{+}$ $=f(x_{\Omega}+h\,(\mathbf{u}+\mathbf{v})),$
and the backward offsets
$f_{\mathbf{u}}^{-}$ $=f(x_{\Omega}-h\,\mathbf{u}),$ $\quad f_{\mathbf{v}}^{-}$ $=f(x_{\Omega}-h\,\mathbf{v}),$
$f_{\mathbf{uv}}^{-}$ $=f(x_{\Omega}-h\,(\mathbf{u}+\mathbf{v})).$
Define the one-sided mixed differences and their symmetric average
$D_{\mathbf{uv}}^{(+)}$ $=\frac{f_{\mathbf{uv}}^{+}-f_{\mathbf{u}}^{+}-f_{\mathbf{v}}^{+}+f_{00}}{h^{2}},$
$D_{\mathbf{uv}}^{(-)}$ $=\frac{f_{\mathbf{uv}}^{-}-f_{\mathbf{u}}^{-}-f_{\mathbf{v}}^{-}+f_{00}}{h^{2}},$
$D_{\mathbf{uv}}^{(\mathrm{c})}$ $=\tfrac{1}{2}\!\left(D_{\mathbf{uv}}^{(+)}+D_{\mathbf{uv}}^{(-)}\right).$
Taylor justification (order and bias). Assume $f\in C^{3}$ near $x_{\Omega}$. Using
$f(x_{\Omega}+h\mathbf{a})=f_{00}+h\,\nabla f^{\top}\mathbf{a}+\tfrac{1}{2}h^{2}\,\mathbf{a}^{\top}H_{f}\,\mathbf{a}+\mathcal{O}(h^{3}),$
with $\mathbf{a}\in\{\mathbf{u},\mathbf{v},\mathbf{u}+\mathbf{v}\}$, we obtain
$f_{\mathbf{uv}}^{+}-f_{\mathbf{u}}^{+}-f_{\mathbf{v}}^{+}+f_{00}$ $=\tfrac{1}{2}h^{2}\!\left[(\mathbf{u}+\mathbf{v})^{\top}H_{f}(\mathbf{u}+\mathbf{v})-\mathbf{u}^{\top}H_{f}\mathbf{u}-\mathbf{v}^{\top}H_{f}\mathbf{v}\right]$
$\quad+\mathcal{O}(h^{3})$
$=h^{2}\,\mathbf{u}^{\top}H_{f}\,\mathbf{v}+\mathcal{O}(h^{3}),$
hence $D_{\mathbf{uv}}^{(+)}=\mathbf{u}^{\top}H_{f}\,\mathbf{v}+\mathcal{O}(h)$. Replacing $h\!\to\!-h$ gives $D_{\mathbf{uv}}^{(-)}=\mathbf{u}^{\top}H_{f}\,\mathbf{v}+\mathcal{O}(h)$. Averaging cancels odd powers:
(9) $D_{\mathbf{uv}}^{(\mathrm{c})}=\mathbf{u}^{\!\top}H_{f}(x_{\Omega})\,\mathbf{v}\;+\;\mathcal{O}(h^{2}).$
@@ 4.1. Off-Diagonal Weingarten Loss via Taylor Expansion
$S_{12}(x_{\Omega})=\frac{\mathbf{u}^{\!\top}H_{f}(x_{\Omega})\,\mathbf{v}}{\|\nabla f(x_{\Omega})\|}.$
Using the finite-difference approximation gives the off-diagonal Weingarten loss
(10) $\bigl{|}S_{12}(x_{\Omega})\bigr{|}\;\approx\;\frac{\bigl{|}D_{\mathbf{uv}}^{(\mathrm{c})}(x_{\Omega})\bigr{|}}{\|\nabla f(x_{\Omega})\|}\;+\;\mathcal{O}(h^{2}),$
@@
(11) $S_{12}^{2}(x_{\Omega})\;\approx\;\frac{\bigl{(}D_{\mathbf{uv}}^{(\mathrm{c})}(x_{\Omega})\bigr{)}^{2}}{\|\nabla f(x_{\Omega})\|^{2}}\;+\;\mathcal{O}(h^{2}).$
In practice we may omit the division by $\|\nabla f(x_{\Omega})\|$ and minimize $\lvert D_{\mathbf{uv}}^{(\mathrm{c})}\rvert$ directly: because the Eikonal loss enforces $\|\nabla f\|\!\approx\!1$ in the shell, both definitions coincide up to a small scale factor. This symmetric stencil converts the curvature-gap loss into a first-order computation.
4.2. Off-Diagonal Weingarten Loss via Auto-Diff
For applications that favor numerical precision we exploit reverse-mode auto-diff to obtain a Hessian–vector product
@@ 4.2. Off-Diagonal Weingarten Loss via Auto-Diff
The first backward sweep delivers $\nabla f$; a second sweep differentiates the scalar inner product $(\nabla f)\!\cdot\!\mathbf{v}$ with respect to the input coordinates, giving the exact mixed second derivative without materializing the full $3{\times}3$ Hessian. This doubles the backward sweeps for the Weingarten loss samples and avoids the prohibitive memory footprint of full Hessian assembly.
This Hessian–vector product therefore requires only two reverse-mode sweeps and stores no matrix entries, so its cost grows linearly with the number of shell samples and is independent of the ambient dimension. Dividing by $\|\nabla f(x_{\Omega})\|$ provides the exact $S_{12}(x_{\Omega})$ value used in the off-diagonal loss. Although more expensive than the first-order finite-difference loss, this second-order route offers higher accuracy and serves as our reference implementation for ablation studies.
@@ 5. Experiments
In this section, we first describe the implementation details and evaluation metrics used in our validation. We then present a comprehensive comparison of our proposed off-diagonal Weingarten based loss functions (ODW-AD and ODW-FD) against three baseline methods-DiGS (Ben-Shabat et al., 2021), NSH (Wang et al., 2023), and the state-of-the-art NeurCADRecon  (Dong et al., 2024)—on two subsets of the ABC dataset (Koch et al., 2019). Finally, we analyze convergence behavior and runtime performance.
@@ 5.1. Experimental Setup
All experiments are conducted on two 100‐example subsets of the ABC dataset (Koch et al., 2019). The first subset (1 MB set) is a pseudo‐random selection of 100 CAD models (each $\approx$ 1 MB in file size), while the second subset (5 MB set) comprises 100 more complex human‐chosen models (between 3.5 MB and 9.5 MB, in average $\approx$ 5 MB) selected for their clean topology and well‐defined features. For each mesh, we uniformly sample a total of 30,000 points to form the input point cloud. This preprocessing strategy ensures that all methods receive identical input distributions. After that, at each iteration 20,000 points are randomly chosen from the 30,000 points. Additional 20,000 non-manifold points are drawn via spatial sampling within the mesh’s bounding volume.
@@
We compare two variants of our off-diagonal loss: ODW-AD (Auto-Diff) and ODW-FD (Finite-Diff) against three baselines: DiGS(Ben-Shabat et al., 2021), Neural-Singular-Hessian (NSH) (Wang et al., 2023) and NeurCADRecon (NCR) (Dong et al., 2024). All methods have been implemented within a unified comparison framework to eliminate discrepancies arising from differing network architectures or training pipelines. Each method employs a SIREN MLP architecture (Sitzmann et al., 2020) with four hidden layers of 256 units and sine activations. All models use a standard SIREN initialization, except DiGS, which uses the MFGI initialization. Training is performed using the Adam optimizer (Kingma and Ba, 2017) with a learning rate of $5\times 10^{-5}$ for up to 10,000 iterations. Early stopping is triggered if the Chamfer-Distance metric plateaus for 1500 consecutive iterations. All other hyperparameters, such as weight coefficients for curvature or normal losses, and other method specific parameters follow the settings from the original implementations.
@@ 5.2. Evaluation Metrics
We evaluate reconstruction accuracy using three standard metrics. Chamfer Distance (CD), scaled by $10^{3}$, which measures the similarity between two surfaces; lower values indicate better accuracy. F1 Score (F1) is the harmonic mean of precision and recall, computed using a distance threshold of $5\times 10^{-3}$ between the predicted and ground-truth point sets. The results are scaled by $10^{2}$, with higher values indicating a better overlap. Normal Consistency (NC) captures the mean cosine similarity between predicted and ground-truth normals, scaled by $10^{2}$; higher values denote better alignment. For each of these metrics, we report the mean and standard deviation across all 100 shapes in each data subset. Additionally, for the 1MB subset, we include additional efficiency measures: GPU VRAM usage (in GB), Time per iteration (in milliseconds), and Best Iteration achieved during evaluation—the iteration count at which the best CD metric is reached. We utilize the time measurements and best iteration count to calculate the convergence time.
@@ 5.3. Qualitative Results
Beyond aggregate scores, the spatial distribution of reconstruction error is inspected in Figure 3 on a Hausdorff-distance heat map. Our method markedly reduces peak deviations and suppresses local error spikes compared with the baselines. Other qualitative results are shown in Figure 1 and in the collection in Figure 10. We also compare our method with other state-of-the-art reconstruction methods, depicted in Figure 4, which shows our approach achieves comparable reconstruction quality with NeurCADRecon (Dong et al., 2024).
@@ 5.4. Input Data Experiments
Input Size NC $\uparrow$ $\mathrm{CD}_{L_{1}}$ F1 $\uparrow$
@@
To evaluate robustness to missing regions, a circular patch on the surface is excised from each input cloud while keeping the remaining samples unchanged. As reported in Table 2 and Figure 6, the Weingarten loss-based reconstruction degrades gracefully: Chamfer distance increases by only $64\,\%$ and normal consistency drops by $0.7\,\%$, yet global topology and salient features are preserved, however, a visible bump remains.
@@
Our method is specifically designed for CAD data, where the Weingarten loss discourages the reconstructed surface to be locally warped and twisted. To understand how this inductive bias affects performance on general, non-CAD geometry, we evaluate our model on the standard Armadillo shape, which contains rich organic, but also high frequency, details.
As shown in Figure 7 and Table 3, our method produces smooth and topologically consistent reconstructions, comparable to those generated by NeurCADRecon (Dong et al., 2024), but at approximately half the training time. Compared to DiGS (Ben-Shabat et al., 2021) and Neural-Singular-Hessian (NSH) (Wang et al., 2023), both our method and NeurCADRecon yield superior global shape recovery, exhibiting fewer holes and more coherent topology. However, we observe that fine surface details are partially smoothed out—an expected consequence of the curvature regularization, which favors larger planar or uniformly curved regions.